OpenStudy (anonymous):

The perimeter of a rectangle is 50 cm and the length is 4 cm longer than the width. Write and solve the system of equations using the substitution method. A. width = 10 cm, length= 14 cm B. width = 10 1/2 cm, length= 14 1/2 cm C. width = 11 cm, length= 15 cm D. width = 23 cm, length= 27 cm

OpenStudy (anonymous):

l + w =50 l = w +4 w+4+w=50 w= (50-4)/2 =23 l = 23+4= 27